OpenStudy (anonymous):

For the following system, if you isolated x in the second equation to use the Substitution Method, what expression would you substitute into the first equation? 2x + y = 8 -x - 3y = -12 3y + 12 -3y + 12 3y - 12 -3y - 12

OpenStudy (anonymous):

-x -3y = -12 +3y +3y -x = 3y -12 -(-x) = -(3y -12) x = -3y +12